A reaction vessel contains 0.0200 M N2 , 0.0600 M H2 , and 0.0100M NH3. With the aid of calculation, predict whether ammonia,NH3 , Will be formed or dissociated when the mixture goes to equilibrium at 400 °C .Oc for the reaction is 0.500 at 400 °C.
1
Expert's answer
2021-10-18T03:09:33-0400

N2 (g) + 3H2 (g) 2NH3 (g)



0.06×400/0.5×400=24/200=0.12